The members of a flying club plan to share equally the cost of $200,000 airplane. The members want to find five more people to join the club so that the cost per person will decrease by $2,000. How many members are currently in the club?
Answer by checkley77(12844)

200,000/x=200,000/(x+5)+2,000
200,000/x-200,000/(x+5)=2,000
[200,000*(x+5)-200,000x]/x(x+5)=2,000
[200,000x+1,000,000)-200,000x)/(x^2+5x)=2,000
2,000(x^2+5x)=1,000,000
2,000x^2+10,000x-1,000,000=0
2,000(x^2+5x-500)=0
2,000(x+25)(x-20)=0
x-20=0
x=20 the current number of members.
Proof:
200,000/20=200,000/25+2,000
10,000=8,000+2,000
10,000=10,000